# Modular reduction of complex representations of finite reductive groups

### v1

@@ Abstract.
Given a complex representation of a finite group, Brauer and Nesbitt defined in 1941 its reduction mod $p$, obtaining a representation over the algebraic closure of $\mathbb{F}_{p}$. In 2021, Lusztig studied the characters obtained by reducing mod $p$ an irreducible unipotent representation of a finite reductive group over $\mathbb{F}_{p}$. He gave a conjectural formula for this character as a linear combination of terms which had no explicit definition and were only known in some small-rank examples. In this paper we provide an explicit formula for these terms and prove Lusztig’s conjecture, giving a formula for the reduction mod $p$ of any unipotent representation of $G(\mathbb{F}_{q})$ for $q$ a power of $p$. We also propose a conjecture linking this construction to the full exceptional collection in the derived category of coherent sheaves on a partial flag variety constructed recently by Samokhin and van der Kallen.
@@ 1. Introduction
Given a complex representation $\rho$ of a finite group $\Gamma$, one can consider the reduction mod $p$ of this representation, an associated representation $\underline{\rho}$ over the field $\overline{\mathbb{F}}_{p}$. This operation was defined by Brauer in Nesbitt in [BN]. In [L], Lusztig studied this operation in the case where $\Gamma=G(\mathbb{F}_{p})$, a finite reductive group over a field of characteristic $p$. He presented a conjecture toward the goal of giving an explicit formula for the character of a representation obtained by applying this operation to any irreducible unipotent representation of $G(\mathbb{F}_{p})$.
@@
No formula for $M_{w}$ was known in general, but these $M_{w}$ were further conjectured to satisfy some desirable properties, most importantly being in a sense “independent of $p$,” which we explain below. In Section 2 of [L], Lusztig computed the elements $M_{w}$ in Types $A_{1},A_{2},B_{2},G_{2}$, and $A_{3}$.
In the present paper, our main construction is an explicit definition of $M_{w}$. Let $T$ be a torus for $G$ split over $\mathbb{F}_{p}$ and let $W$ be the Weyl group with longest element $w_{0}$. We study the module $\mathbb{C}[T]$ over $\mathbb{C}[T]^{W}$ and construct a system of generators $\{f_{w}\}_{w\in W}$ with a simple and explicit definition which satisfy many useful properties. There is a $\mathbb{C}[T]^{W}$-linear endomorphism $[q]^{*}$ of $\mathbb{C}[T]$ given on this basis by $[q]^{*}f_{w}(t)=f_{w}(t^{q})$. We consider also a standard $\mathbb{C}[T]^{W}$-linear pairing $\langle$,$\rangle$ , ⟩ on $\mathbb{C}[T]$ given by the Weyl character formula, obtaining the following result.
@@
We recall the definition of a Duflo involution in Section 4.1. In Types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, and $A_{3}$ considered in [L], the $M_{w}$ we define agree with the ones computed in loc. cit. To any $w\in W$, we associate a sum of fundamental weights $\omega_{I(w)}$ determined by its right descent set, which we explain in Section 3.1. In [L], Lusztig also suggests some properties that one might expect the elements $M_{w}$ to satisfy.
Conjecture 1.3 ([L]).
@@
    For any $w\in\mathcal{J}$, $M_{w}$ is a linear combination of characters of Weyl modules $V_{\lambda}$ for $\lambda$ very close to $(q-1)\omega_{I(w)}$.
@@
Despite this, in addition to providing counterexamples we give a natural explanation for the failure of these properties, having to do with a discrepancy between the dual basis $\{f_{w}^{*}\}_{w\in W}$ and a certain pseudo-dual basis $f^{w}$ which we also define. This discrepancy only appears in types beyond those for which Lusztig explicitly computed the elements $M_{w}$ in [L]. We view this (in combination with some uniqueness results we examine in Section 6) as compelling evidence for why the properties in Conjecture 1.3 were satisfied in Lusztig’s examples but not in general.
In conclusion, our definition of the elements $M_{w}$ and subsequent proof of the formula (1) provide an explicit expansion of the Brauer reduction $\underline{\rho}$ of an irreducible unipotent representation $\rho$ of $G(\mathbb{F}_{q})$ as a linear combination of Weyl modules restricted to $G(\mathbb{F}_{q})$.
The elements $M_{w}$ arise in the present approach as a computational tool for evaluating the trace of operators induced by $\tilde{\phi}\colon\mathcal{O}(T)\to\mathcal{O}(T)$, which can be thought of as an $\mathcal{O}(T)^{W}$-linear approximation to the Frobenius endomorphism of the ring $\mathcal{O}(T)$. It would be very interesting to find a natural categorification of the operator $\tilde{\phi}$ and/or of the elements $M_{w}$, perhaps finding an object of motivic nature that produces both a representation of $G(\mathbb{F}_{q})$ with characteristic zero coefficients and a module over the algebraic group. The conjecture on the relation to the exceptional collection of [SVK] stated in Section 7 is a step in this direction; we plan to return to this question in the future.
Acknowledgments. The first author wishes to thank Bao Le Hung for drawing his attention to [L]. We would like to thank Alexander Samokhin and Wilberd van der Kallen for helpful communications and for generously sharing their code related to the paper [SVK]. The research of D.K. was partially supported by ERC grant 101142781, R.B. was partly supported by the NSF grant DMS-2101507. The collaboration of R.B. and D.K. was partly supported by the US-Israel BSF grant 2020189.
@@ 2.1. Preliminaries
Let $\mu$ be the group of roots of unity in $\mathbb{C}^{\times}$. We fix once and for all a surjective homomorphism $\psi\colon\mu\to\mathsf{k}^{\times}$ whose kernel is the subset of $\mu$ consisting of elements whose order is a power of $p$. For any finite group $\Gamma$ and a representation $\rho$ of $\Gamma$ over $\mathbb{C}$, Brauer and Nesbitt in [BN] defined an associated representation $\underline{\rho}$ of $\Gamma$ over $\mathsf{k}$, which we call the reduction of $\rho$ mod $p$. The representation $\underline{\rho}$ is characterized by the property that for any $g\in\Gamma$, the eigenvalues of $g$ on $\underline{\rho}$ are obtained by applying $\psi$ to the eigenvalues of $g$ on $\rho$.
Following [L, Section 0.1], we let $\mathcal{R}G(\mathbb{F}_{q})$ (resp. $\mathcal{R}_{p}G(\mathbb{F}_{q})$) be the Grothendieck group of finite-dimensional representations of $G(\mathbb{F}_{q})$ over $\mathbb{C}$ (resp. over $\mathsf{k}$), with $\mathcal{R}^{+}G(\mathbb{F}_{q})$ (resp. $\mathcal{R}_{p}^{+}G(\mathbb{F}_{q})$) the submonoid consisting of classes of actual representations. We use $\mathcal{R}_{p}G$ to denote representations of the algebraic group $G$ over $\mathsf{k}$.
For any $\lambda\in X^{*}(T)^{+}$, the set of dominant weights, we let $V_{\lambda}\in\mathcal{R}_{p}^{+}G(\mathbb{F}_{q})$ be the character of the Weyl module corresponding to $\lambda$, which is given by the Weyl character formula. For a $p$-restricted weight $\lambda$, we let $L_{\lambda}$ be the corresponding irreducible representation of $G(\mathbb{F}_{q})$ as explained in [L]. When $\lambda=n_{1}\varpi_{1}+\dots+n_{r}\varpi_{r}$ for fundamental weights $\varpi_{k}$, we set $V_{n_{1},\dots,n_{r}}=V_{\lambda}$, $L_{n_{1},\dots,n_{r}}=L_{\lambda}$.
@@
Let $[q]\colon T\to T$ be given by $z\mapsto z^{q}$, let $\pi\colon T\to T/W$ be the projection and $\overline{[q]}\colon T/W\to T/W$ be the unique morphism such that $\pi\circ[q]=\overline{[q]}\circ\pi$. Let $(T/W)_{q}$ be the fixed point scheme of $\overline{[q]}$, and note that it is reduced. Let $\iota_{q}$ be the inclusion map $\iota_{q}\colon(T/W)_{q}\hookrightarrow T/W$. The sheaf $\pi_{*}(\mathcal{O}_{T})$ carries an action of $W$ while $\iota_{q}^{*}\pi_{*}(\mathcal{O}_{T})$ carries an automorphism $\phi$ defined by $\phi(f)=[q]^{*}f$. For convenience, we write $A_{q}=\mathcal{O}((T/W)_{q})$.
We will also make frequent use of the canonical basis $\{C_{w}\}_{w\in W}$ introduced by Kazhdan-Lusztig in [KL] and the dual canonical basis $\{C_{w}^{\prime}\}_{w\in W}$ for $\mathbb{Z}[W]$. We recall that for any $s\in S$,
@@
The basis $C_{w}$ was used in [KL] to define a partition of the set $W$ into left-, right-, or two-sided Kazhdan-Lusztig cells. Any such set of Kazhdan-Lusztig cells admits a partial order $\leq$. For any two-sided cell $c$, we write
@@ 2.2. Unipotent representations and almost-characters
For any $\chi\in\mathrm{Irr}(W)$, there is a corresponding complex irreducible unipotent representation $U_{\chi}$ of $G(\mathbb{F}_{q})$. The representations of the form $U_{\chi}$ for some $\chi\in\mathrm{Irr}(W)$ are called irreducible unipotent principal series representations. One can also associate to $\chi$ an element $V_{\chi}\in\mathcal{R}G(\mathbb{F}_{q})$ called an almost character, the virtual unipotent representation of $G(\mathbb{F}_{q})$ over $\mathbb{C}$ whose character is the trace of the characteristic function of the corresponding Frobenius-invariant character sheaf on $G$, as introduced in [LBook].
In [LBook, Section 12.1], Lusztig also introduced the non-abelian Fourier transform for representations of $G(\mathbb{F}_{q})$. It can be thought of as an endomorphism $\mathrm{FT}$ of $\mathcal{R}G(\mathbb{F}_{q})$ satisfying the property that for any $\chi\in\mathrm{Irr}(W)$, it exchanges $V_{\chi}$ with the character of $U_{\chi}$.
@@ 2.3. The Brauer character
This is an equivalent form of formula (1) in [J, Section 3.2] for $w=1$. ∎
@@
To define $\tilde{\phi}$, and therefore $\tilde{V}_{\chi}$, we will proceed as follows. Given a basis $f_{w}$ for the free $A$-module $N$, one gets a uniquely defined $A$-module endomorphism such that $\tilde{\phi}(f_{w})=[q]^{*}f_{w}$. If the $\mathbb{C}$-span of $f_{w}$ is $W$-invariant then $\tilde{\phi}$ commutes with the action of $W$, so Corollary 2.3 applies. Notice that $\tilde{\phi}$ only depends on the $\mathbb{C}$-span of $f_{w}$.
@@ 3.1. Preliminaries
For $i\in\Pi$, let $\omega_{i}$ be the corresponding fundamental weight. For any subset $I\subseteq\Pi$, we write $\overline{I}$ for its complement in $\Pi$. Let $\omega_{I}=\sum_{i\in I}\omega_{i}$. We can consider the parabolic subgroup $W_{I}\subset W$ generated by $\{s_{i}\}_{i\in I}$. Recall that for any $w$,
@@
In [St], a basis $\{E_{w}\}_{w\in W}$ for $N=\mathbb{C}[T]$ over $A=\mathbb{C}[T]^{W}$ was constructed, which we refer to as the Steinberg basis. Explicitly, for any $w\in W$, we have
$E_{w}=w\exp(\omega_{I(w)}).$
@@ 3.2. The Kazhdan-Lusztig-Steinberg basis
This means once we define a basis satisfying the conditions in Lemma 3.1, if we set $\tilde{V}_{\chi}=\frac{1}{\dim\chi}\mathrm{tr}(\tilde{\phi},[\mathrm{gr}_{w}(N):\chi])$ for the unique equivalence class of $w$ for which $[\mathrm{gr}_{w}(N):\chi]\neq 0$, then the equation
@@
Let (,)(\ ,\ )( , ) be the standard pairing on $\mathbb{C}[W]$ given by
@@
$f_{w}$ $=\frac{1}{|W_{I(w)}|}C_{w}^{\prime}(\exp(\omega_{\overline{I(w)}})),$
$f^{w}$ $=\frac{1}{|W_{\overline{I(w)}}|}C^{w}(\exp(\omega_{I(w)})).$
Note that for any $w\in W$, $f_{w}\in N_{\mathbb{Z}}$. This is because one can write $C_{w}^{\prime}=C_{w}^{\prime\prime}\cdot a_{I(w)}$ for some $C_{w}^{\prime\prime}\in\mathbb{Z}[W]$, where $a_{I(w)}$ is the idempotent of the trivial representation. The claim then follows from the fact that $a_{I(w)}\exp(\omega_{\overline{I(w)}})=|W_{I(w)}|\exp(\omega_{\overline{I(w)}})$.
@@
One can check that if $h(v)<h(w)$, then $x\omega_{\overline{I(w)}}+y\omega_{I(v)}$ is a singular weight (i.e. has nontrivial stabilizer under $W$) for any $x,y\in W$ (note that $\overline{I(w)}=I(w_{0}w)$). This also holds if $h(v)=h(w)$ and $I(v)\neq I(w)$. It then follows from the Weyl character formula that $\langle f_{w},f^{v}\rangle=0$ if $h(v)\leq h(w)$ and $I(v)\neq I(w)$.
@@
$\langle w\exp(\omega_{I(w)}),\exp(\omega_{\overline{I(w)}})\rangle$ =$=\begin{cases}0&w\not\in W_{I(w)}W_{\overline{I(w)}}\\ (-1)^{x}&w=xy,x\in W_{I(w)},y\in W_{\overline{I(w)}},\end{cases}$
@@
First note that $\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}=\mathbb{Z}[W]^{<c}N$. Indeed, $\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}\subset\mathbb{Z}[W]^{\leq c}N$ by definition of $f_{y}$. Choosing a direct complement $N=N^{\prime}\oplus\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}$, we see that if $\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}\subsetneq\mathbb{Z}[W]^{\leq c}N$, then $\mathbb{Z}[W]^{\leq c}N$ has rank strictly bigger than |$|\{y\in W\mid y\leq c\}|$. This implies that the same is true of $\mathbb{Z}[W]^{\leq c}N/A_{+}$ as a subspace of $N/A_{+}\cong H^{\bullet}(G/B)$ (where $A_{+}\subset A$ denotes the subspace generated by terms of positive degree) but this is a contradiction by the well-known structure of $H^{\bullet}(G/B)$ as a $W$-module.
For any $v,w$ in a left cell $c_{L}\subset c$, we have $I(v)=I(w)$ by [LUnequal, Section 8.6]. For a fixed weight $\lambda$, the map $w\mapsto w\exp(\lambda)$ is clearly $W$-equivariant. It follows that the map $\mathbb{Z}[W]^{c_{L}}\to N^{c_{L}}_{\mathbb{Z}}$ given by $C^{\prime}_{w}\mapsto f_{w}$, is $W$-equivariant. Here $\mathbb{Z}[W]^{c_{L}}\subset\mathbb{Z}[W]^{c}$ is spanned by $C^{\prime}_{w}$, $w\in c_{L}$. Further, $\mathbb{Z}[W]^{c}=\bigoplus_{c_{L}\subset c}\mathbb{Z}[W]^{c_{L}}$. It is then an isomorphism by part (1). ∎

### v2

@@ Abstract.
The main result describes the Brauer-Nesbitt reduction of unipotent representations of a finite group of Lie type, expressing it as an explicit linear combination of the restriction of Weyl modules from the algebraic group to the group of $\mathbb{F}_{q}$ points. This partly confirms Lusztig’s conjecture (2021), which was the main source of motivation for this work.
The explicit virtual representations of the algebraic group come from a certain endomorphism of the space ${\mathbb{Z}}[T]$ of regular functions on the torus which approximates pullback under Frobenius and is linear over the ring ${\mathbb{Z}}[T]^{W}$ of $W$-invariant functions. This endomorphism is constructed from a new basis for ${\mathbb{Z}}[T]$ over ${\mathbb{Z}}[T]^{W}$ which we call the Kazhdan-Lusztig-Steinberg basis.
We compare this basis to the canonical basis appearing in the study of modular representations of the algebraic group and the related noncommutative Springer resolution. This leads to canonically defined objects in the derived category of $G$-modules representing the above virtual representations and to a geometric interpretation for the resulting lift of the principal series representation $\overline{\mathbb{F}_{q}}[G/P(\mathbb{F}_{q})]$ to a virtual representation of the algebraic group, which comes from a decomposition of diagonal in the equivariant Grothendieck group of the partial flag variety.
@@ 1. Introduction
Given a complex representation $\rho$ of a finite group $\Gamma$, one can consider the reduction mod $p$ of this representation, an associated representation $\underline{\rho}$ over the field $\overline{\mathbb{F}}_{p}$. This operation was defined by Brauer and Nesbitt in [undefd]. In [undefp], Lusztig studied this operation in the case where $\Gamma=G(\mathbb{F}_{p})$, a finite reductive group over a field of characteristic $p$. He presented a conjecture toward the goal of giving an explicit formula for the character of a representation obtained by applying this operation to any irreducible unipotent representation of $G(\mathbb{F}_{p})$.
@@
No specific formula for $M_{w}$ was proposed in general, but these $M_{w}$ were further conjectured to satisfy some desirable properties, most importantly being in a sense “independent of pp,” which we explain below. In Section 2 of [undefp], Lusztig computed the elements $M_{w}$ in Types $A_{1},A_{2},B_{2},G_{2}$, and $A_{3}$.
In the present paper, our main construction is an explicit definition of $M_{w}$. Let $T$ be a torus for $G$ split over $\mathbb{F}_{p}$ and let $W$ be the Weyl group with longest element $w_{0}$. We study the module $\mathbb{C}[T]$ over $\mathbb{C}[T]^{W}$ and construct a system of generators $\{f_{w}\}_{w\in W}$ with a simple and explicit definition which satisfy many useful properties, which we call the Kazhdan-Lusztig-Steinberg basis, as it shares some properties both with the canonical bases and with the Steinberg basis of ${\mathbb{Z}}[W]$ from [undefw]. There is a $\mathbb{C}[T]^{W}$-linear endomorphism $q]^{*}$ of $\mathbb{C}[T]$ given on this basis by $q]^{*}f_{w}(t)=f_{w}(t^{q})$. We consider also a standard $\mathbb{C}[T]^{W}$-linear pairing $\langle,\rangle$ on $\mathbb{C}[T]$ given by the Weyl character formula, obtaining the following result.
@@
We recall the definition of a Duflo involution in Section 4.1. In Types $A_{1}$, $A_{2}$, $B_{2}$, $G_{2}$, and $A_{3}$ considered in [undefp], the $M_{w}$ we define agree with the ones computed in loc. cit. To any $w\in W$, we associate a sum of fundamental weights $\varpi_{I(w)}$ determined by its right descent set, which we explain in Section 3.1. In [undefp], Lusztig also suggests some properties that one might expect the elements $M_{w}$ to satisfy.
Conjecture 1.3 ([undefp]).
@@
    For any $w\in\mathcal{J}$, $M_{w}$ is a linear combination of characters of Weyl modules $V_{\lambda}$ for $\lambda$ very close to $(q-1)\varpi_{I(w)}$.
@@
Despite this, in addition to providing counterexamples we give a natural explanation for the failure of these properties, having to do with a discrepancy between the dual basis $\{f_{w}^{*}\}_{w\in W}$ and a certain pseudo-dual basis $f^{w}$ which we also define. This discrepancy only appears in types beyond those for which Lusztig explicitly computed the elements $M_{w}$ in [undefp]. We view this (in combination with some uniqueness results we examine in Section 6) as compelling evidence for why the properties in Conjecture 1.3 were satisfied in Lusztig’s examples but not in general.
Thus our definition of the elements $M_{w}$ and subsequent proof of the formula (1) provide an explicit expansion of the Brauer reduction $\underline{\rho}$ of an irreducible unipotent representation $\rho$ of $G(\mathbb{F}_{q})$ as a linear combination of Weyl modules restricted to $G(\mathbb{F}_{q})$.
The elements $M_{w}$ arise in the above argument as a computational tool for evaluating the trace of operators induced by $\tilde{\phi}\colon\mathcal{O}(T)\to\mathcal{O}(T)$, which can be thought of as an $\mathcal{O}(T)^{W}$-linear approximation to the Frobenius endomorphism of the ring $\mathcal{O}(T)$. In Section 7 we lift $M_{w}$ to naturally defined objects in the derived category of $G$-modules, and describe a geometric way to lift the principal series $G(\mathbb{F}_{q})$-module $\overline{\mathbb{F}_{q}}[(G/P)(\mathbb{F}_{q})]$ to a virtual representation of the algebraic group $G$ compatible with the above lift of an arbitrary unipotent representation.
A general observation (see Lemma 7.1) links such a lift for the module $\overline{\mathbb{F}_{q}}[X(\mathbb{F}_{q})]$, where $X$ is a projective variety over $\mathbb{F}_{q}$ equipped with a $G$-action, to decomposition of diagonal in the equivariant $K$-group $K^{G}(X^{2})$. We show that noncommutative Springer resolution yields such a decomposition for the partial flag variety, and check compatibility with the above, based on the relation of the noncommutative Springer resolution with the canonical bases.
It is tempting to speculate that the virtual representations $M_{w}$ should admit a categorification in a still stronger sense, perhaps finding an object of motivic nature that produces both a representation of $G(\mathbb{F}_{q})$ with characteristic zero coefficients and a module over the algebraic group.
To facilitate computations, we have made available an implementation of the constructions in the present paper in a public repository [undefu]. Readers can use it to reproduce all of the computations in the present paper and to experiment further with the constructions introduced here.
Acknowledgments. The first author wishes to thank Bao Le Hung for drawing his attention to [undefp]. We would like to thank Alexander Samokhin and Wilberd van der Kallen for helpful communications and for generously sharing their code related to the paper [undefv]. The research of D.K. was partially supported by ERC grant 101142781, R.B. was partly supported by the NSF grant DMS-2101507. The collaboration of R.B., M.F. and D.K. was partly supported by the US-Israel BSF grant 3013008329.
@@ 2.1. Preliminaries
Let $\mu$ be the group of roots of unity in $\mathbb{C}^{\times}$. We fix once and for all a surjective homomorphism $\psi\colon\mu\to\mathsf{k}^{\times}$ whose kernel is the subset of $\mu$ consisting of elements whose order is a power of $p$. For any finite group $\Gamma$ and a representation $\rho$ of $\Gamma$ over $\mathbb{C}$, Brauer and Nesbitt in [undefd] defined an associated representation $\underline{\rho}$ of $\Gamma$ over $\mathsf{k}$, which we call the reduction of $\rho$ mod $p$. The representation $\underline{\rho}$ is characterized by the property that for any $g\in\Gamma$, the eigenvalues of $g$ on $\underline{\rho}$ are obtained by applying $\psi$ to the eigenvalues of $g$ on $\rho$.
Following [undefp, Section 0.1], we let $\mathcal{R}G(\mathbb{F}_{q})$ (resp. $\mathcal{R}_{p}G(\mathbb{F}_{q})$) be the Grothendieck group of finite-dimensional representations of $G(\mathbb{F}_{q})$ over $\mathbb{C}$ (resp. over $\mathsf{k}$), with $\mathcal{R}^{+}G(\mathbb{F}_{q})$ (resp. $\mathcal{R}_{p}^{+}G(\mathbb{F}_{q})$) the submonoid consisting of classes of actual representations. We use $\mathcal{R}_{p}G$ to denote representations of the algebraic group $G$ over $\mathsf{k}$.
For any $\lambda\in X^{*}(T)^{+}$, the set of dominant weights, we let $V_{\lambda}\in\mathcal{R}_{p}^{+}G(\mathbb{F}_{q})$ be the character of the Weyl module corresponding to $\lambda$, which is given by the Weyl character formula. For a $p$-restricted weight $\lambda$, we let $L_{\lambda}$ be the corresponding irreducible representation of $G(\mathbb{F}_{q})$ as explained in [undefp]. When $\lambda=n_{1}\varpi_{1}+\dots+n_{r}\varpi_{r}$ for fundamental weights $\varpi_{k}$, we set $V_{n_{1},\dots,n_{r}}=V_{\lambda}$, $L_{n_{1},\dots,n_{r}}=L_{\lambda}$.
@@
Let $q]\colon T\to T$ be given by $z\mapsto z^{q}$, let $\pi\colon T\to T/W$ be the projection and $\overline{[q]}\colon T/W\to T/W$ be the unique morphism such that $\pi\circ[q]=\overline{[q]}\circ\pi$. Let $(T/W)_{q}$ be the fixed point scheme of $\overline{[q]}$, and note that it is reduced. Let $\iota_{q}$ be the inclusion map $\iota_{q}\colon(T/W)_{q}\hookrightarrow T/W$. The sheaf $\pi_{*}(\mathcal{O}_{T})$ carries an action of $W$ while $\iota_{q}^{*}\pi_{*}(\mathcal{O}_{T})$ carries an automorphism $\phi$ defined by $\phi(f)=[q]^{*}f$. For convenience, we write $A_{q}=\mathcal{O}((T/W)_{q})$.
We will also make frequent use of the canonical basis $\{C_{w}\}_{w\in W}$ introduced by Kazhdan-Lusztig in [undefk] and the dual canonical basis $\{C_{w}^{\prime}\}_{w\in W}$ for $\mathbb{Z}[W]$. We recall that for any $s\in S$,
@@
The basis $C_{w}$ was used in [undefk] to define a partition of the set $W$ into left-, right-, or two-sided Kazhdan-Lusztig cells. Any such set of Kazhdan-Lusztig cells admits a partial order $\leq$. For any two-sided cell $c$, we write
@@ 2.2. Unipotent representations and almost-characters
For any $\chi\in\mathrm{Irr}(W)$, there is a corresponding complex irreducible unipotent representation $U_{\chi}$ of $G(\mathbb{F}_{q})$. The representations of the form $U_{\chi}$ for some $\chi\in\mathrm{Irr}(W)$ are called irreducible unipotent principal series representations. One can also associate to $\chi$ an element $V_{\chi}\in\mathcal{R}G(\mathbb{F}_{q})$ called an almost character, the virtual unipotent representation of $G(\mathbb{F}_{q})$ over $\mathbb{C}$ whose character is the trace of the characteristic function of the corresponding Frobenius-invariant character sheaf on $G$, as introduced in [undefs].
In [undefs, Section 12.1], Lusztig also introduced the non-abelian Fourier transform for representations of $G(\mathbb{F}_{q})$. It can be thought of as an endomorphism $\mathrm{FT}$ of $\mathcal{R}G(\mathbb{F}_{q})$ satisfying the property that for any $\chi\in\mathrm{Irr}(W)$, it exchanges $V_{\chi}$ with the character of $U_{\chi}$.
@@ 2.3. The Brauer character
This is an equivalent form of formula (1) in [undefi, Section 3.2] for $w=1$. ∎
@@
To define $\tilde{\phi}$, and therefore $\tilde{V}_{\chi}$, we will proceed as follows. Given a basis $f_{w}$ for the free $A$-module NN, one gets a uniquely defined $A$-module endomorphism such that $\tilde{\phi}(f_{w})=[q]^{*}f_{w}$. If the $\mathbb{C}$-span of $f_{w}$ is $W$-invariant then $\tilde{\phi}$ commutes with the action of $W$, so Corollary 2.3 applies. Notice that $\tilde{\phi}$ only depends on the $\mathbb{C}$-span of $f_{w}$.
@@ 3.1. Preliminaries
For $i\in\Pi$, let $\varpi_{i}$ be the corresponding fundamental weight. For any subset $I\subseteq\Pi$, we write $\overline{I}$ for its complement in $\Pi$. Let $\varpi_{I}=\sum_{i\in I}\varpi_{i}$. We can consider the parabolic subgroup $W_{I}\subset W$ generated by $\{s_{i}\}_{i\in I}$. Recall that for any $w$,
@@
In [undefw], a basis $\{e_{w}\}_{w\in W}$ for $N=\mathbb{C}[T]$ over $A=\mathbb{C}[T]^{W}$ was constructed, which we refer to as the Steinberg basis. Explicitly, for any $w\in W$, we have
$e_{w}=w\exp(\varpi_{I(w)}).$
@@ 3.2. The Kazhdan-Lusztig-Steinberg basis
This means once we define a basis satisfying the conditions in Lemma 3.1, if we set $\tilde{V}_{\chi}=\frac{1}{\dim\chi}\mathrm{tr}(\tilde{\phi},[\mathrm{gr}_{w}(N):\chi])$ for the unique equivalence class of $w$ for which $\mathrm{gr}_{w}(N):\chi]\neq 0$, then the equation
@@
Let (,)(\ ,\ ) be the standard pairing on $\mathbb{C}[W]$ given by
@@
$f_{w}$ $=\frac{1}{|W_{I(w)}|}C_{w}^{\prime}(\exp(\varpi_{\overline{I(w)}})),$
$f^{w}$ $=\frac{1}{|W_{\overline{I(w)}}|}C^{w}(\exp(\varpi_{I(w)})).$
Note that for any $w\in W$, $f_{w}\in N_{\mathbb{Z}}$. This is because one can write $C_{w}^{\prime}=C_{w}^{\prime\prime}\cdot(|W_{I(w)}|a_{I(w)})$ for some $C_{w}^{\prime\prime}\in\mathbb{Z}[W]$, where $a_{I(w)}=\frac{1}{|W_{I(w)}|}\sum_{w\in W_{I(w)}}w$ is the idempotent of the trivial representation of the parabolic subgroup $W_{I(w)}$. The claim then follows from the fact that $a_{I(w)}\exp(\varpi_{\overline{I(w)}})=\exp(\varpi_{\overline{I(w)}})$.
@@
One can check that if $h(v)<h(w)$, then $x\varpi_{\overline{I(w)}}+y\varpi_{I(v)}$ is a singular weight (i.e. has nontrivial stabilizer under $W$) for any $x,y\in W$ (note that $\overline{I(w)}=I(w_{0}w)$). This also holds if $h(v)=h(w)$ and $I(v)\neq I(w)$. It then follows from the Weyl character formula that $\langle f_{w},f^{v}\rangle=0$ if $h(v)\leq h(w)$ and $I(v)\neq I(w)$.
@@
$\langle w\exp(\varpi_{I(w)}),\exp(\varpi_{\overline{I(w)}})\rangle$ $=\begin{cases}0&w\not\in W_{I(w)}W_{\overline{I(w)}}\\ (-1)^{x}&w=xy,x\in W_{I(w)},y\in W_{\overline{I(w)}},\end{cases}$
@@
First note that $\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}=\mathbb{Z}[W]^{\leq c}N$. Indeed, $\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}\subset\mathbb{Z}[W]^{\leq c}N$ by definition of $f_{y}$. Choosing a direct complement $N=N^{\prime}\oplus\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}$, we see that if $\bigoplus_{y\leq c}A_{\mathbb{Z}}f_{y}\subsetneq\mathbb{Z}[W]^{\leq c}N$, then $\mathbb{Z}[W]^{\leq c}N$ has rank strictly bigger than $|\{y\in W\mid y\leq c\}|$. This implies that the same is true of $\mathbb{Z}[W]^{\leq c}N/A_{+}$ as a subspace of $N/A_{+}\cong H^{\bullet}(G/B)$ (where $A_{+}\subset A$ denotes the subspace generated by terms of positive degree) but this is a contradiction by the well-known structure of $H^{\bullet}(G/B)$ as a $W$-module.
For any $v,w$ in a left cell $c_{L}\subset c$, we have $I(v)=I(w)$ by [undefn, Section 8.6]. For a fixed weight $\lambda$, the map $w\mapsto w\exp(\lambda)$ is clearly $W$-equivariant. It follows that the map $\mathbb{Z}[W]^{c_{L}}\to N^{c_{L}}_{\mathbb{Z}}$ given by $C^{\prime}_{w}\mapsto f_{w}$, is $W$-equivariant. Here $\mathbb{Z}[W]^{c_{L}}\subset\mathbb{Z}[W]^{c}$ is spanned by $C^{\prime}_{w}$, $w\in c_{L}$. Further, $\mathbb{Z}[W]^{c}=\bigoplus_{c_{L}\subset c}\mathbb{Z}[W]^{c_{L}}$. It is then an isomorphism by part (1). ∎